OpenStudy (anonymous):

If an object traveling at 5 m/s [N] accelerates 8 m/s [NW], then what is the resultant velocity?

OpenStudy (anonymous):

This question is not possible. We need the amount of time the object accelerated for or some indication of a final velocity.

OpenStudy (anonymous):

I was thinking the same thing lol sorry friend posted me this question and I thought I was doing something wrong

OpenStudy (anonymous):

Well. It increases 8 ms/s per second. So if you by chance get the time you can easily find out the result.

OpenStudy (anonymous):

This can easily be solved with the information given.

OpenStudy (anonymous):

the displacement is 10.5 m/s since the object is going 5 m/s already and it changes direction 45 degrees so that makes its new velocity 2.5 m/s but then it accelerates 8 m/s faster than before making the new velocity 10.5 m/s

OpenStudy (anonymous):

I may be mistaken but I don't think that roravets answer is correct. If you give each velocity a vector representation and add them you get two sides of a triangle with an angle in between. This means all you have to do is solve for the third side of the triangle with the cosine rule. a^2 = b^2 + c^2 - 2abCos(A) or... answer^2 = 64 + 25 - 80*Cos(135) The final answer I get is a = 12.065 = 12